# Don’t Retrieve, Navigate: Distilling Enterprise Knowledge into Navigable Agent Skills for QA and RAG

### 29 Apr 2026 v2

@@ Abstract
1 Introduction
Figure 1: Retrieve vs. Navigate. Traditional RAG passively feeds fixed passages to the LLM. Corpus2Skill distills the corpus into a navigable skill hierarchy that the agent actively explores, backtracks, and drills into to locate evidence.
Enterprise knowledge bases contain thousands of heterogeneous documents: support articles, product guides, policy documents, and FAQs (Gao et al., 2024). Large language models alone cannot answer questions about these corpora reliably: they lack access to private, frequently updated content and are prone to hallucinating plausible but incorrect details (Lewis et al., 2020; Gao et al., 2024). Retrieval-Augmented Generation (RAG; Lewis et al., 2020) was introduced to solve these problems by grounding LLM responses in retrieved evidence. By embedding the query, fetching similar passages from a vector index, and conditioning the model’s answer on the retrieved context, RAG enables LLMs to leverage enterprise-specific knowledge they were never trained on, providing factual, up-to-date answers while reducing hallucination. This pipeline has become the dominant paradigm for enterprise QA, customer support, and compliance workflows (Gao et al., 2024; Singh et al., 2025). Yet it presents a structural limitation for more complex queries: the LLM never sees the forest for the trees.
Consider a support agent faced with the query “How do I switch my business from sole proprietorship to LLC on Wix?” This query spans multiple topics: account types, payment provider configuration, and legal entity management, each covered in different sections of the knowledge base. A flat retrieval system returns the top-$k$ passages closest to the query embedding, which are likely to match on surface-level terms (“sole proprietorship,” “LLC”) but miss the operationally critical article explaining that account types cannot be changed directly and require contacting support. The LLM receives these fragments but has no understanding of how the corpus is organized, what topic areas exist beyond the retrieved set, or whether the best-matching documents were even in the retrieval window. It cannot reason about where else to look, whether its evidence is complete, or how the retrieved passages relate to the broader knowledge structure. In short, conventional RAG treats the LLM as a passive consumer of search results, leaving the model’s planning and reasoning capabilities largely unused.
Agentic RAG systems (Yao et al., 2023; Schick et al., 2023) partially address this by allowing the LLM to issue multiple search queries iteratively. But even with tool access, the agent operates without a map: it must guess productive search terms, has no way to survey the corpus structure, and cannot systematically narrow its focus from broad topics to specific documents. Each search query is a shot in the dark, guided only by the agent’s prior knowledge and whatever fragments previous queries returned. For the business-type query above, an agentic retrieval system would need to independently discover that the answer involves both the payments ecosystem and account verification policies — topics whose connection is not apparent from keyword similarity alone.
Hierarchical approaches such as RAPTOR (Sarthi et al., 2024) and GraphRAG (Edge et al., 2024) organize documents into trees or graphs with LLM-generated summaries at internal nodes. These summaries capture thematic structure that flat embeddings miss. However, these systems still rely on embedding-based similarity search at query time: RAPTOR indexes all tree nodes in a flat vector store, while GraphRAG requires a graph database and entity extraction pipeline. The hierarchy enriches the retrieval candidates but remains invisible to the LLM at query time; the model still receives search results without understanding the organizational structure that produced them. More recent work pushes further: StructRAG (Li et al., 2025) restructures retrieved information into tables or graphs at inference time, and BookRAG (Wang et al., 2025) builds hierarchical indices with agent-based planners. Yet these systems still invoke embedding-based search at query time and do not give the agent direct access to the organizational structure itself.
Meanwhile, a parallel line of work on LLM agent skills (Wang et al., 2023; Jiang et al., 2026) has shown that agents perform complex tasks more effectively when given structured, reusable knowledge modules that they can selectively load and follow. Agent platforms now support filesystem-based skill packages [^1] https://docs.anthropic.com/en/docs/agents-and-tools/agent-skills that the agent discovers via metadata and reads on demand, a mechanism known as progressive disclosure (Section 2.3). This suggests an alternative to retrieval: rather than searching for passages, distill the corpus into a navigable skill hierarchy and let the agent explore it.
We propose exactly this: make the hierarchy itself the primary interface for the LLM. Rather than searching over a hierarchy, the agent navigates through it. This is the core idea behind Corpus2Skill (Figure 1): a one-time offline compilation phase transforms a document corpus into a navigable skill tree, and at serve time, the LLM agent browses this tree to find information, much like a human navigating a well-organized file system or table of contents.
We evaluate Corpus2Skill on WixQA (Cohen et al., 2025), an enterprise customer-support benchmark designed for evaluating RAG systems over real-world knowledge bases. Corpus2Skill achieves the highest quality across all metrics, outperforming flat retrieval, RAPTOR, and a multi-turn agentic RAG system, while requiring no embedding index or vector database at serve time. We complement this anchor result with a generalization study on nine RAGBench (Friel et al., 2024) subsets reformulated as retrieval benchmarks: Corpus2Skill attains the highest macro-average F1 over the full ten-dataset benchmark, and the per-dataset dispersion isolates the regime where navigation is the right primitive (single-domain, atomic-document corpora) versus the regime where flat retrieval remains preferable (open-domain abstracts, very long extractive documents). These results suggest a complementary design point for knowledge-grounded systems: invest computation at compile time to make the corpus navigable, rather than solely at query time to make it searchable.
We summarize our contributions as follows.
  1. 1.
    We identify corpus navigability as an underexplored dimension in current RAG systems: LLMs lack a structured overview of the knowledge base, which limits their ability to systematically locate the best evidence.
  2. 2.
    We propose Corpus2Skill, a compile-then-navigate framework: an offline pipeline transforms a raw corpus into a hierarchical skill directory via iterative clustering and LLM-based summarization, and at serve time an LLM agent navigates this hierarchy through file browsing and document lookup tools, requiring no embedding index or vector database at serve time.
  3. 3.
    We evaluate on WixQA (Cohen et al., 2025) against five baselines, conduct ablation studies on hierarchy structure, exploration budget, and serving model choice, and run a 9-subset generalization study on RAGBench (Friel et al., 2024) that characterizes which corpus shapes Corpus2Skill helps and which it does not.
2 Background: Skills in LLM Agents
2.1 From Tools to Skills
Modern LLM agents interact with external systems through tools: atomic, stateless functions such as web search or database queries (Schick et al., 2023; Yao et al., 2023). Tools give an agent the ability to act but do not tell it how to approach a complex task. Agent skills fill this gap: reusable modules that package procedural knowledge or multi-step workflows into a form the agent can load and follow (Jiang et al., 2026). Unlike a tool, a skill shapes reasoning over multiple steps, persisting across tool calls to guide the agent’s decisions throughout a task. This distinction is increasingly important as agents tackle longer-horizon tasks where choosing the right sequence of tool calls requires domain-specific strategy, not just tool availability.
Formally, Jiang et al. (2026) define a skill as a tuple $S=(C,\pi,T,R)$: an applicability condition $C$, an execution policy $\pi$, a termination condition $T$, and a reusable interface $R$. This distinguishes skills from one-off plans (no reusable interface), episodic memories (no execution policy), and atomic tools (no multi-step structure).
2.2 Skill Representations
Skills take several forms in practice:
  • Code-as-skill. Voyager (Wang et al., 2023) stores successful action sequences as executable functions in a growing library.
  • Natural-language playbooks. SkillX (Wang et al., 2026) and EvoSkill (Alzubi et al., 2026) distill agent experience into structured textual instructions.
  • Filesystem-based packages. A directory containing a structured instruction file (SKILL.md) alongside optional scripts and data assets. This is the pattern adopted by recent agent platforms and the one Corpus2Skill builds on.
Code-based skills offer precision but require execution environments; natural-language skills are portable and interpretable at the cost of ambiguity. Filesystem-based packages balance the two by combining structured text with auxiliary resources that the agent can access on demand.
2.3 Progressive Disclosure
Loading all skill content at startup would overwhelm the context window; loading nothing would leave the agent unable to choose. The filesystem pattern resolves this via progressive disclosure (Shneiderman, 1996):
  1. 1.
    Metadata (at startup): skill name and one-line description (<<100 tokens each).
  2. 2.
    Instructions (on demand): the full SKILL.md is read when the agent selects a skill.
  3. 3.
    Resources (as needed): supporting files remain on disk, consuming zero tokens until accessed.
The agent thus maintains awareness of many skills while paying the token cost only for those actually used.
2.4 From Procedural Skills to Informational Skills
In prior work, skills encode how to perform a task: workflows, decision trees, or code patterns. Corpus2Skill repurposes the same mechanism to encode what a corpus contains. Each skill directory represents a topical partition, with SKILL.md and INDEX.md files that summarize its contents at increasing levels of detail. The agent follows the same progressive disclosure workflow (description $\to$ SKILL.md $\to$ INDEX.md $\to$ document) but navigates an information hierarchy rather than executing procedural instructions. This reinterpretation—from procedural to informational skills—is the central design insight of Corpus2Skill.
3 Corpus2Skill Framework
Corpus2Skill operates in two phases: a one-time offline compile phase that transforms a raw document corpus into a browsable skill tree, and a per-query online serve phase where an LLM agent navigates this tree with full visibility into the corpus structure. Figure 2 illustrates the full architecture.
3.1 Problem Setup
Let $\mathcal{D}=\{d_{1},\dots,d_{N}\}$ be a corpus of $N$ documents. Given a natural-language query $q$, the task is to produce an answer $a$ grounded in a subset $\mathcal{D}_{q}\subseteq\mathcal{D}$ of relevant documents. Standard RAG systems select $\mathcal{D}_{q}$ via a retrieval function $f(q,\mathcal{D})\to\mathcal{D}_{q}$ that operates as a black box from the LLM’s perspective: the model receives $\mathcal{D}_{q}$ but has no visibility into what $\mathcal{D}$ contains or why $\mathcal{D}_{q}$ was chosen.
Corpus2Skill replaces this opaque retrieval step with agent-driven navigation. The compile phase constructs a hierarchical skill forest $\mathcal{S}=\{s_{1},\dots,s_{K}\}$ where each skill $s_{k}$ is a navigable directory tree of summaries covering a topical partition of $\mathcal{D}$. Following the agentic skills formalism of Jiang et al. (2026), each skill can be viewed as a tuple $(C,\pi,T,R)$: the applicability condition $C$ is the skill’s summary (the agent reads it to decide relevance); the execution policy $\pi$ is the navigation workflow (descend through summaries, then retrieve documents); the termination condition $T$ is the agent’s judgment that sufficient evidence has been gathered; and the reusable interface $R$ is the standardized file structure (SKILL.md, INDEX.md) and the get_document tool. At serve time, the agent selects and navigates skills autonomously, making the document selection process transparent and steerable.
Figure 2: Corpus2Skill system architecture. The compile phase (top) embeds documents, builds a multi-level cluster hierarchy with LLM summarization at each level, and materializes the result as a forest of navigable skills. Root clusters map to skill directories (SKILL.md), sub-clusters map to group directories (INDEX.md), and leaf documents are stored in a separate document store. The serve phase (bottom) uses an LLM agent with pre-compiled skills and two tools: file navigation (code_execution) and document lookup (get_document).
3.2 Compile Phase
The compilation pipeline consists of four stages (implementation details and full prompts in Appendices A and C).
Document loading and embedding.
The compiler accepts directories containing standard document formats (.md, .txt, .json, .jsonl). Each document is assigned a content-hash-based unique identifier, ensuring deterministic, collision-free IDs that are independent of the source system’s naming scheme. Documents are truncated to a maximum character length and embedded using a sentence embedding model (Reimers and Gurevych, 2019), producing a dense vector for each document. These embeddings capture semantic similarity between documents and serve as the foundation for the subsequent clustering step.
Iterative hierarchical clustering.
We build a multi-level hierarchy through an iterative bottom-up process controlled by two parameters: the branching ratio $p$ (target number of children per cluster) and the maximum number of top-level clusters $K$. At each level, the current set of $n$ embeddings is partitioned into $\lceil n/p\rceil$ groups via K-Means (Lloyd, 1982) on L2-normalized vectors. Each resulting cluster is then summarized by an LLM that reads representative member texts and produces a concise description of the cluster’s topical coverage, the types of questions the documents answer, and key terms or features mentioned. These summaries are re-embedded using the same sentence embedding model, and the resulting embeddings become the input to the next round of K-Means. This process repeats until the number of remaining clusters falls below $K$, yielding a set of root clusters that form the top level of the hierarchy.
This iterative embed-cluster-summarize loop is inspired by the recursive tree construction in RAPTOR (Sarthi et al., 2024), but adapted for our navigability goal. We use K-Means with hard assignments so that each document maps to exactly one path in the directory tree, and we materialize the hierarchy as browsable files rather than indexing it in a flat vector store. These two choices, deterministic partitioning and filesystem materialization, are what enable agent-driven navigation at serve time. Small clusters below a minimum size are merged into the nearest centroid to avoid orphan groups.
Labeling.
After the clustering hierarchy stabilizes, the LLM generates a short (2 to 5 word) filesystem-safe label for each non-leaf node. These labels serve a dual purpose: they provide human-readable directory names for inspection (e.g., wix-commerce-monetization) and, more importantly, they give the agent semantically meaningful entry points for topic routing. When the agent reads a SKILL.md listing subgroups such as “wix-payments-ecosystem” or “domain-dns-management,” the labels alone often suffice to identify the relevant branch without reading full summaries.
Skill tree construction.
The hierarchy is materialized into a filesystem structure designed for agent navigation:
  • Root clusters become top-level skill directories, each containing a SKILL.md with the cluster summary and a listing of child groups.
  • Sub-clusters become subdirectories with INDEX.md files that list either further subgroups or leaf document identifiers alongside brief summaries of each document.
  • Full document text is stored externally in a documents.json file, referenced by document ID from within the navigation files.
This separation between navigation metadata and document content is a deliberate architectural choice. Navigation files (SKILL.md, INDEX.md) remain small (typically under 2 KB each), containing only summaries and document identifiers. This keeps the agent’s per-step token cost low: reading an INDEX.md to survey 20 documents costs far fewer tokens than reading the documents themselves. Full document content is accessed only when the agent has narrowed its search, via a dedicated get_document tool call. The result is a forest of $K$ independent skill trees, each representing a topical domain within the corpus. The forest structure allows the agent to reason about topic relevance at the coarsest level before committing to a branch. [^2] Anthropic’s current Skills API imposes limits of 8 skills per request, 200 files and 30 MB per skill. These constraints informed our default parameters but are not fundamental to the framework design.
Skill file design.
Each SKILL.md and INDEX.md follows a uniform template with YAML frontmatter (name, description, level, num_documents) that the Skills API reads for progressive disclosure. Compilation prompts produce routing-oriented summaries: the LLM identifies the topic area, question types, and key terms, so that each file functions as a decision point for the navigating agent. Each file also embeds actionable next-step instructions (e.g., “Use get_document with the doc_id”), so the agent always knows which tool to invoke next.
Figure 3 shows abridged examples of both file types. The SKILL.md (top) gives the agent a single-glance overview of all subgroups within a skill; the leaf INDEX.md (bottom) lists individual document IDs with concise titles for targeted retrieval. Unabridged versions at all three hierarchy levels are in Appendix B.
SKILL.md (root level, abridged) ---
name: skill-01-wix-commerce-monetization
description: >
Full Wix commerce and monetization
ecosystem, from storefronts to payments.
level: 3
num_documents: 1513
## Contents
### Sub-groups (directories)
- group-00-wix-restaurants/ (159 docs):
Wix Restaurants orders, menus...
- group-01-wix-payments-ecosystem/ (87):
Payment processing, account mgmt...
- group-02-wix-pos-systems/ (110 docs)
[...16 subgroups total]
INDEX.md (leaf level, verbatim) ---
name: L1-C298
description: >
Fraud prevention and payment security
within Wix Payments.
level: 1 num_documents: 5
### Documents (5 items)
- 5̀6ae76:̀ Best Practices for Accepting
Card Payments
- 4̀48d79:̀ About the ‘Liability Shift’
- b̀b73d9:̀ Declined Payment Due to
Suspected Fraud
- c̀3b01a:̀ About Suspicious Payments
- d̀22b8c:̀ Double Charge on Statement
Figure 3: Abridged SKILL.md (top) and leaf INDEX.md (bottom) from the WixQA compilation. The SKILL.md provides a routing overview of 1,513 documents; the INDEX.md lists individual document IDs with titles for targeted retrieval.

### 15 May 2026 v3

@@ 5 Experiments
BM25 0.345 0.830 0.472 0.831 0.406 0.549 0.005 1.00 0.007
Dense 0.364 0.833 0.537 0.889 0.448 0.659 0.000 1.00 0.007
Hybrid 0.361 0.832 0.524 0.851 0.416 0.637 0.005 1.00 0.007
RAPTOR 0.398 0.839 0.704 0.909 0.618 0.659 0.005 1.00 0.012
Agentic 0.378 0.843 0.705 0.528 0.452 0.649 0.500 5.06 0.082
Corpus2Skill 0.456 0.862 0.767 0.859 0.708 0.829 0.045 2.38 0.153
Table 2: Generalization on 10 RAGBench subsets (200 queries each, or all available) over the expanded retrieval pool. Datasets grouped by the structural pattern that determines whether navigation helps. Per-dataset best in bold; quality cells green, \$/q red.
Dataset Method Quality Metrics Cost
[…]
(a) Single-domain, atomic-document corpora — Corpus2Skill wins
Emanual BM25 0.412 0.855 0.661 0.874 0.556 0.586 0.000 1.00 0.006
Dense 0.421 0.855 0.691 0.894 0.650 0.668 0.000 1.00 0.006
Hybrid 0.424 0.854 0.721 0.895 0.647 0.644 0.000 1.00 0.006
Agentic 0.455 0.871 0.841 0.696 0.611 0.633 0.220 3.39 0.046
Corpus2Skill 0.497 0.898 0.817 0.835 0.768 0.832 0.091 2.57 0.060
DelucionQA BM25 0.412 0.858 0.730 0.946 0.681 0.664 0.005 1.00 0.006
Dense 0.440 0.862 0.821 0.969 0.750 0.753 0.000 1.00 0.006
Hybrid 0.429 0.861 0.775 0.957 0.714 0.733 0.000 1.00 0.006
Agentic 0.438 0.874 0.841 0.719 0.711 0.725 0.206 2.91 0.039
Corpus2Skill 0.478 0.893 0.821 0.921 0.824 0.872 0.022 2.64 0.110
TechQA BM25 0.377 0.840 0.537 0.781 0.295 0.432 0.040 1.00 0.009
Dense 0.382 0.840 0.560 0.774 0.297 0.499 0.050 1.00 0.009
Hybrid 0.383 0.840 0.562 0.779 0.267 0.466 0.040 1.00 0.009
Agentic 0.309 0.836 0.605 0.414 0.271 0.476 0.635 6.33 0.151
Corpus2Skill 0.396 0.854 0.709 0.672 0.496 0.649 0.280 3.75 0.507
ExpertQA BM25 0.312 0.825 0.486 0.913 0.439 0.408 0.005 1.00 0.007
Dense 0.330 0.827 0.573 0.909 0.529 0.507 0.010 1.00 0.007
Hybrid 0.326 0.827 0.552 0.898 0.478 0.463 0.015 1.00 0.007
Agentic 0.288 0.838 0.709 0.517 0.442 0.457 0.508 4.39 0.085
Corpus2Skill 0.356 0.854 0.764 0.666 0.565 0.589 0.225 3.88 0.183
ClapNQ BM25 0.325 0.849 0.735 0.940 0.766 0.406 0.000 1.00 0.005
Dense 0.333 0.850 0.774 0.934 0.788 0.420 0.005 1.00 0.005
Hybrid 0.333 0.850 0.756 0.938 0.782 0.411 0.005 1.00 0.005
Agentic 0.308 0.855 0.756 0.736 0.710 0.421 0.161 2.52 0.027
Corpus2Skill 0.468 0.893 0.924 0.911 0.937 0.878 0.015 2.03 0.080
(b) Heterogeneous or numerical-table corpora — flat retrieval competitive, Corpus2Skill matches
CovidQA BM25 0.204 0.785 0.546 0.905 0.542 0.517 0.050 1.00 0.006
Dense 0.215 0.804 0.635 0.927 0.634 0.602 0.030 1.00 0.006
Hybrid 0.216 0.795 0.613 0.924 0.608 0.602 0.035 1.00 0.006
Agentic 0.206 0.849 0.651 0.612 0.549 0.617 0.363 4.13 0.058
Corpus2Skill 0.208 0.847 0.456 0.735 0.405 0.426 0.255 6.85 0.436
FinQA BM25 0.359 0.847 0.340 0.940 0.358 0.413 0.015 1.00 0.005
Dense 0.400 0.852 0.520 0.924 0.514 0.479 0.020 1.00 0.006
Hybrid 0.389 0.853 0.440 0.936 0.438 0.481 0.005 1.00 0.006
Agentic 0.368 0.861 0.576 0.618 0.512 0.506 0.420 4.36 0.076
Corpus2Skill 0.392 0.867 0.574 0.770 0.550 0.652 0.175 4.51 0.495
(c) Open-domain factoid or homogeneous-tabular corpora — Corpus2Skill loses
HAGRID BM25 0.320 0.862 0.815 0.967 0.825 0.613 0.010 1.00 0.005
Dense 0.322 0.862 0.855 0.974 0.881 0.627 0.000 1.00 0.005
Hybrid 0.323 0.862 0.831 0.970 0.859 0.641 0.005 1.00 0.005
Agentic 0.356 0.879 0.826 0.881 0.829 0.623 0.071 2.29 0.022
Corpus2Skill 0.188 0.838 0.358 0.787 0.220 0.200 0.260 1.05 0.021
TatQA BM25 0.307 0.847 0.468 0.965 0.471 0.489 0.005 1.00 0.005
Dense 0.312 0.846 0.527 0.935 0.508 0.498 0.020 1.00 0.006
Hybrid 0.315 0.848 0.536 0.951 0.522 0.537 0.010 1.00 0.005
Agentic 0.355 0.866 0.612 0.638 0.537 0.510 0.390 3.02 0.045
Corpus2Skill 0.240 0.847 0.388 0.706 0.367 0.499 0.285 6.92 1.475
CUAD BM25 0.314 0.824 0.277 0.962 0.181 0.477 0.000 1.00 0.007
Dense 0.304 0.819 0.279 0.963 0.167 0.742 0.005 1.00 0.008
Hybrid 0.308 0.821 0.277 0.945 0.195 0.685 0.000 1.00 0.007
Agentic 0.249 0.819 0.270 0.402 0.177 0.706 0.695 4.27 0.080
Corpus2Skill 0.272 0.828 0.237 0.643 0.217 0.439 0.430 5.80 0.502
Table 3: Ablations on WixQA (200 queries): (a) cluster structure, (b) agent exploration budget, (c) serving LLM. Per-panel best in bold.
[…]
(a) Cluster structure (pp/skills/depth)
Narrow (5/2/5) 0.462 0.864 0.758 0.690 0.080 2.38 0.133
Default (10/6/3) 0.456 0.862 0.767 0.708 0.045 2.38 0.153
Wide (20/4/3) 0.382 0.847 0.552 0.414 0.245 3.69 0.565
(b) Agent exploration budget (max rounds)
5 rounds 0.455 0.863 0.774 0.702 0.065 2.35 0.145
10 rounds 0.459 0.863 0.754 0.693 0.065 2.36 0.142
20 rounds 0.456 0.862 0.767 0.708 0.045 2.38 0.153
(c) Serving LLM (default tree, max_turns=20{=}20)
Sonnet 4.6 0.456 0.862 0.767 0.708 0.045 2.38 0.153
Haiku 4.5 0.420 0.856 0.671 0.722 0.180 2.92 0.093
5.1 Setup
Dataset and compilation.
We evaluate on WixQA (Cohen et al., 2025), a customer-support QA benchmark with 6,221 support articles and 200 expert-written test queries with gold answers and gold article IDs. We compile WixQA with branching ratio p=10p{=}10 and maximum top-level clusters K=10K{=}10, producing a 3-level hierarchy with 6 top-level skills and 665 navigation files. Compilation takes 19.5 minutes on an RTX 4090 server.
Baselines.
All methods use the same answer LLM and the same evaluation protocol. We compare five baselines spanning three retrieval paradigms: BM25 (sparse keyword retrieval (Robertson and Zaragoza, 2009)), Dense (Qwen3-Embedding-0.6B (Zhang et al., 2025) with a FAISS index (Johnson et al., 2019)), Hybrid (RRF (Cormack et al., 2009) of BM25 and Dense), RAPTOR (hierarchical tree retrieval (Sarthi et al., 2024) with UMAP+GMM+BIC clustering), and Agentic (an LLM agent with iterative access to BM25, Dense, and Hybrid tools). The first three are single-shot (top-5 passages, one LLM call). RAPTOR uses collapsed tree retrieval, and Agentic is allowed up to 10 rounds. Full implementation in Appendix˜H.
Metrics.
[…]
5.2 Main Results
Answer quality.
Corpus2Skill achieves the highest score on every answer-quality metric and on both retrieval-coverage metrics (Table˜1). Token F1 reaches 0.456 (a 21% relative gain over Agentic 0.378 and 25% over Dense 0.364). BERTScore-F1 is 0.862, ahead of Agentic (0.843), RAPTOR (0.839), and the flat retrievers (0.830–0.833). Factuality 0.767 exceeds the next-best methods Agentic (0.705) and RAPTOR (0.704) by 6 absolute points. CtxR/CtxP lead by clear margins (0.708/0.829 vs. 0.618/0.659 for RAPTOR and 0.452/0.649 for Agentic). Both hierarchical methods outperform the flat baselines, but Corpus2Skill extends RAPTOR’s gains by a further 9% on Factuality and 15% on CtxR: cluster summaries surface thematic connections flat embeddings miss, letting the agent match queries to relevant document groups rather than passages.
Faithfulness and hallucination.
[…]
Cost.
[…]
5.3 Generalization to RAGBench
To probe how Corpus2Skill behaves outside its target domain, we run the same five methods on ten RAGBench (Friel et al., 2024) subsets (CovidQA (Möller et al., 2020), DelucionQA (Sadat et al., 2023), Emanual (Nandy et al., 2021), ExpertQA (Malaviya et al., 2024), TechQA (Castelli et al., 2020), FinQA (Chen et al., 2021), CUAD (Hendrycks et al., 2021), HAGRID (Kamalloo et al., 2023), TatQA (Zhu et al., 2021), ClapNQ (Rosenthal et al., 2025)), sampling 200 test queries per subset (or all available). RAGBench is published as a generator-evaluation benchmark with per-row gold contexts. To convert it into a retrieval benchmark we build the expanded retrieval pool as the union of all train+val+test documents (deduplicated by content hash, pool sizes 221–12,727). This is intentionally harder than the original RAGBench protocol: every test query competes against thousands of distractors.
Corpus navigation is not one-size-fits-all.
The cross-dataset picture is decidedly mixed: Corpus2Skill wins on 7 of 11 datasets, ties on 2, and loses on 3 (Table˜2). The win/loss split is not random but instead tracks the structure of the underlying corpus, and three regimes separate wins from losses. (i) Single-domain, topically coherent corpora where each document covers one feature or topic (WixQA, Emanual, DelucionQA, TechQA, ExpertQA, ClapNQ): Corpus2Skill beats every retrieval baseline by 0.01–0.14 F1, with ClapNQ the most striking case (0.468 vs. Dense 0.333, ++0.135 absolute). (ii) Heterogeneous or numerical-table corpora (CovidQA, FinQA): all methods sit within 0.02 F1, and we treat these as ties. (iii) Open-domain factoid or homogeneous-tabular corpora (HAGRID, TatQA, CUAD): Corpus2Skill loses by 0.04–0.17 F1 because the level-1 cluster summaries collapse onto generic labels (HAGRID’s open-domain mix), describe near-identical units (TatQA’s repeated SEC tables), or partition a single long document across multiple clause types (CUAD).
Grounding and positioning.
[…]
5.4 Ablation Studies
We conduct three ablation studies on WixQA to characterize how compile-time and serve-time design choices affect the quality-cost tradeoff (Table˜3). All variants share the two-tool agent configuration and are directly comparable to Table˜1.
Cluster structure (panel a).
Varying $p$ at K=10K{=}10 produces narrow (p=5p{=}5), default (p=10p{=}10), and wide (p=20p{=}20) trees. Narrow and default land within $\approx$1 F1 point (0.462 vs. 0.456) on every quality metric, and narrow is slightly cheaper (\$0.133 vs. \$0.153). The wide shape collapses to F1 0.382 (0.07-0.07), CtxR 0.414 (0.29-0.29), Hallucination 24.5%, 3.7 turns / \$0.57 per query: once a level-1 cluster must summarize hundreds of topically-mixed documents its summary becomes too generic to route on, reproducing the TatQA failure mode of Section˜5.3 in miniature.
Agent exploration budget (panel b).
Varying max rounds (5/10/20) has minimal impact on quality (within 0.005 F1, avg. 2.4 turns regardless of budget) because the hierarchy is well-organized enough that the agent rarely needs extended exploration. A larger budget mainly helps grounding: Hallucination drops monotonically from 6.5% (budget 5) to 4.5% (budget 20), suggesting that the extra rounds are used to verify rather than re-search.
Serving LLM choice (panel c).
Swapping Sonnet 4.6 for Haiku 4.5 (same compiled tree) cuts cost \$0.153 $\to$ \$0.093 ($39\%$) while retaining 92% of Sonnet’s F1 (0.420) and 87% of its Factuality (0.671). CtxR actually increases (0.708 $\to$ 0.722) because Haiku tends to read one more document before answering. The tradeoff is grounding: Haiku’s Hallucination Rate is $4\times$ Sonnet’s, so Sonnet remains the right choice for compliance-sensitive deployments. The Haiku variant is the cheapest Corpus2Skill configuration we measure and still substantially beats every flat retrieval baseline on F1 by 2–6 absolute points, evidence that the compiled tree, not the navigator’s sophistication, drives retrieval performance. The navigating LLM only has to be good enough to follow signposted instructions.
6 Conclusion
We introduced Corpus2Skill, a framework that compiles a document corpus into a hierarchical skill directory and replaces embedding-based retrieval with agent-driven navigation. On WixQA, the approach improves answer quality and grounding over flat, hierarchical, and agentic baselines. A ten-subset RAGBench generalization study shows that navigation wins on the majority of corpora, with substantial gains on single-domain collections, while flat retrieval remains preferable on open-domain or homogeneous-tabular corpora where top-level clustering collapses. Ablations confirm that the compiled tree, not the serving model, drives performance: even a smaller navigator retains most of the quality at markedly lower cost. Because the output is plain skill directories requiring no vector database, it is platform-agnostic and easy to integrate into existing agent deployments. For corpora with recoverable topical structure, making a corpus navigable complements making it searchable.
Limitations
Cost.
[…]
Routing errors.
Although soft assignment and entity cross-linking (Section˜4.1) let the agent discover cross-cutting content, our failure analysis (Appendix˜G) shows that top-level routing errors still account for the majority of failures. Documents spanning multiple topics can be cross-referenced in secondary branches, but the primary routing decision remains a bottleneck when the query targets the less obvious topic.
Compilation.
The one-time compilation takes $\sim$20 minutes of wall-clock time for WixQA (6,221 documents), negligible for enterprise deployments handling thousands of daily queries. The pipeline is fully automated and deterministic given fixed random seeds, but does not yet support incremental updates: adding documents requires recompilation.
Corpus scope.
Our RAGBench study (Section˜5.3) identifies three corpus shapes where navigation does not help: long extractive documents where answers are verbatim spans (CUAD), homogeneous tabular corpora whose cluster summaries collapse onto near-duplicate labels (TatQA), and open-domain pools where dense retrieval already surfaces sufficient evidence in a single shot (HAGRID). These are structural scope limits rather than implementation issues, and motivate hybrid per-query routing between navigation and retrieval.
Appendix A Compilation Details
Top-level skill labels.
The 6 skills produced by the default configuration (p=10p{=}10, K=10K{=}10) on the WixQA corpus are shown in Table˜4.
Table 4: Top-level skills produced by Corpus2Skill on WixQA.
Skill Label Docs Coverage
commerce-operations 1,682 Commerce, POS, payments
platform-overview 1,770 Editors, CMS, design
platform-ecosystem 1,797 Domains, billing, lifecycle
members-community 256 Members, gated access
bookings-platform 283 Bookings, scheduling
media-display 433 Galleries, storefront
Embedding.
Documents and cluster summaries are embedded using Qwen/Qwen3-Embedding-8B, a 7.6B-parameter sentence encoder producing 4,096-dimensional embeddings. We use the document prompt template with batch size 32 on a single GPU.
Document summary cards.
Before embedding, each document is processed by a lightweight summary-card LLM call (Claude Haiku 4.5, $\leq$250 output tokens, single API call per document, async with a semaphore of 20). The card is a JSON object with fields title, one_line, and phrases (up to four 1–4-word distinctive phrases such as product names or feature flags). The card is concatenated with the (truncated) raw text as a single string and embedded as one vector per document; the card is also retained on the document record and reused later as INDEX.md row metadata, so the per-document summarization cost is amortized across embedding, navigation, and serving.
Cluster summarization.
Cluster summaries are produced by Claude Sonnet 4.6 with direct API calls (no agent loop). Level-1 summaries condense up to 15 document texts (600 characters each); deeper-level summaries condense up to 20 child summaries (300 characters each). Summarization is performed concurrently using async calls with a semaphore of 20. Internal-level clusters are re-embedded by concatenating their summary with a sampled set of member excerpts before passing the result to the next round of K-Means, so each higher-level vector reflects both the abstractive summary and surface-lexical signal.
Soft assignment.
K-Means assignment uses a margin threshold $\tau{=}0.05$ on cosine similarity: items whose top-1 versus top-2 centroid similarity gap is below $\tau$ retain their primary parent and additionally register a secondary parent. Secondary assignments are not duplicated as full documents; the runner-up parent’s INDEX.md lists them as @see stubs pointing back to the canonical location. On the WixQA compile this produces 2,068 secondary stubs spread across 665 navigation files.
LLM verify-and-repartition.
After each level’s clusters are summarized, a verify-and-repartition LLM pass examines all sibling cluster summaries together (in batches of up to 60 clusters per call) and proposes ID-level reassignments where two summaries cover the same topic (confusable) or one summary spans multiple topics (mixed). The pass operates at levels 1 and 2 only and visits each cluster at most once per compile; affected clusters are re-summarized after reassignment. This adds one extra Sonnet call per affected sibling group and a re-summarization call per repartitioned cluster.
Exemplars.
For every non-leaf cluster, the top-3 documents closest (by cosine similarity to the cluster centroid) are surfaced under an ## Example documents section in the rendered SKILL.md/INDEX.md. Exemplars carry the document title and one-line description from the summary card.
Entity extraction and related skills.
Once the cluster hierarchy is stable, a single batched Sonnet call set tags each cluster summary with up to 12 named entities (proper nouns, product names, error codes, file formats) and up to 4 document types. The entity tags are aggregated into a corpus-level entity_index.json that maps each entity to the skill paths that mention it (sorted by mention count). For every top-level skill, the top-3 sibling skills by entity-Jaccard overlap are recorded as ## Related skills, giving the agent explicit cross-jump targets. On the WixQA compile this produces 2,328 distinct entries in entity_index.json.
File counts.
The compilation produces 665 navigation files (SKILL.md + INDEX.md), a documents.json file storing the full text of all 6,221 documents (13 MB), and a entity_index.json mapping 2,328 entities to skill paths. Total output size is 16 MB.